# A Closer Look at AUROC and AUPRC under Class Imbalance

### v3

@@ Results.
Our results demonstrate the impact of the optimization metric on subpopulation disparity. In particular, in Figure 2, we observe a notable disparity introduced when optimizing under the AUPRC metric regardless of the optimization procedure. This is evident in the performance metrics across the high and low prevalence subpopulations, which exhibit significant divergence as the optimization process favors the group with higher prevalence. In the more realistic, random-permutation optimization procedure (Figure 2d), this even results in a decrease in the AUROC for the low prevalence subgroup. In comparison, when optimizing for overall AUROC, the AUROC of both groups increase together. Note that we show the effect of this optimization on the AUPRC metric, which shows very similar trends, in Appendix Figure 4.
@@ 3.2 Real-World Experimental Validation
To demonstrate the generalizability of our finding to the real world, we evaluate fairness gaps induced by AUROC and AUPRC selection on four common datasets in the fairness literature (Zhang et al., 2018; Fabris et al., 2022; Lahoti et al., 2020).
@@
We use the following four tabular binary classification datasets: adult (Asuncion & Newman, 2007), compas (Angwin et al., 2022), lsac (Wightman, 1998), and mimic (Johnson et al., 2016). In each dataset, we consider both sex and race as sensitive attributes. To mimic the setting of our theorems, we balance each dataset by the sensitive attribute during both training and test, by randomly subsampling the majority group. Further details about each dataset, as well as preprocessing steps, can be found in Appendix G.
@@
We train XGBoost models (Chen & Guestrin, 2016) on each dataset. For each task, we iterate over a grid of per-group weights in order to create a diverse set of models that favor different groups. For each setting of task and per-group weight, we conduct a random hyperparameter search (Bergstra & Bengio, 2012) with 50 runs. We evaluate the validation set overall AUROC and AUPRC. We also evaluate the test set AUROC gap and AUPRC gap between groups, where gaps are defined as the value of the metric for the higher prevalence group minus the value for the lower prevalence group. Based on our theorems, our hypothesis is that overall AUPRC should be more positively correlated with the signed AUROC gap than overall AUROC, indicating that it better favors the higher prevalence group, especially when the prevalence ratio between groups is high. To test this hypothesis, we evaluate the Spearman correlation coefficient between these quantities. We repeat this experiment 5 times, with different random data splits, to obtain a 95% confidence interval.
Figure 3: Difference in the Spearman’s $\rho$ between the test-set signed AUROC gap versus the validation set overall AUPRC, and the AUROC gap versus the overall AUROC. Numbers in parentheses are the prevalence ratios between the two groups for the particular attribute, and datasets are sorted by this quantity. Error bars are 95% confidence intervals from five different random data splits.
@@ Results.
In Figure 3, we plot the difference in the Spearman correlation coefficient of the AUROC gap versus the overall AUPRC, and AUROC gap versus overall AUROC. We observe mixed results in datasets with low prevalence ratio. In dataset with higher prevalence ratio, we find that overall AUPRC is more positively correlated with the AUROC gap than overall AUROC, indicating that AUPRC more favors the higher prevalence group. We emphasize that the prevalence ratios observed in these real-world datasets is much lower than the ratio of 5 used in our synthetic experiments, which may account for the mild effect observed. To see raw results from these experiments, see Appendix Figure 7.
Next, in Appendix Figure 8, we plot the difference in the Spearman’s $\rho$ from Figure 3, versus the prevalence gap. We find that there is a statistically significant correlation between the two (Spearman’s $\rho$ = 0.714, p = 0.047). Thus, while our power to detect a prevalence mediated AUPRC bias amplification effect is limited due to the limited prevalence disparities in these datasets, we nonetheless observe a strong positive correlation between the extent of the prevalence mismatch between the low and high prevalence group and the amount that AUPRC favors the high prevalence group over AUROC. In other words, our results show that across these fairness datasets and attributes, as the prevalence disparity grows more extreme, we observe a statistically significant corresponding increase in the extent to which AUPRC introduces algorithmic bias, exactly in accordance with what Theorem 3 suggests.
4 If not for class imbalance, then when should we use AUPRC vs. AUROC?
In Sections 2 and 3, we have shown that AUPRC is not universally superior in cases of class imbalance (and that instead it merely preferentially optimizes high-score regions over low-score regions) and that it also poses serious risks to the model fairness in settings where subgroup prevalences differ. In light of this, how should we revise Claim 1 to reflect when we actually should use AUPRC instead of AUROC or vice versa?
In this section, we explore this question, and provide practical guidance on metric selection. We build on our existing theoretical results and argue that the decision to utilize the AUROC or the AUPRC as the metric for evaluating binary classification models is intricately linked to the specific context in which the model operates, including the relative impact of false positives versus false negatives and the planned model workflow in deployment settings. To demonstrate this, here we return to the four example problems first depicted in Figure 1:
@@
For model evaluations conducted outside of specific deployment contexts, where the differential costs of errors are undefined, the necessity for a metric that impartially values improvements across the entire model output space becomes paramount. As shown in Figure 1a, in this setting, as it is not known in advance where samples of interest will live in the output space nor are particular cost ratios known, correcting any model mistake should be prioritized equally to any other. This scenario inherently advantages AUROC, attributing to its capacity to uniformly account for every correction, thereby offering a comprehensive assessment of model performance irrespective of decision thresholds.
@@
In applications where the consequences of false negatives are especially grave, such as in the early screening for critical illnesses like cancer (refer to Figure 1c), the primary focus of the model in use will be to ensure that as few positive samples are missed as possible. This equates to prioritizing model recall. In such a scenario, the most important mistakes to correct are actually those that occur at lower score thresholds, because high-score mistakes will not change which positive samples are missed in deployment settings as chosen thresholds are likely to be low. This behavior is the inverse of what AUPRC prioritizes, demonstrating that in such situations, AUROC should be preferred over AUPRC. This choice will better ensure a reduction in false negatives, enhancing the model’s proficiency in detecting all possible disease instances.
For ethical resource distribution among large populations, use AUROC:
When faced with the challenge of ethically allocating scarce resources across a broad population, necessitating equitable benefit distribution among subgroups (illustrated in Figure 1d), it is crucial to avoid prioritizing model improvements that selectively favor one subpopulation. As AUPRC will target high-score regions selectively, it risks unduely favoring high-prevalence subpopulations, as shown in Theorem 3 and Figures 2 and 3. Even though in this resource distribution problem, high-score regions are selectively important compared to low-score regions, the fact that in this problem, we must prioritize across all subpopulations equally means that AUPRC’s global preference is untenable as it could induce bias. Thus, as shown in Figure 1d, we should prefer AUROC in this scenario as it will ensure uniform preference across groups, even though it prioritizes both low and high-score regions within each group.
@@
5 If AUPRC is not better than AUROC under class imbalance, why did we think it was?
Claim 1, which states that “AUPRC is better than AUROC in cases of class imbalance” is widespread in the literature. Via both a manual literature search and an automated search of over 1.5M arXiv papers (see Appendix H for methodology), we observed 128 publications making this claim. We analyzed these papers to answer the following two key questions: (1) how has this incorrect claim become so widespread in ML and (2) has this claim led to inappropriate usage of evaluation metrics in high-impact settings?
5.1 Finding 1: This claim is often made without attribution or with inappropriate attribution.
This claim is frequently stated without any citation.
Among the 128 papers we discovered referencing this claim, 31 did so with no associated citation (Liu et al., 2023; Randl et al., 2023; Tusfiqur et al., 2022; Piermarini et al., 2023; Zhang & Bondell, 2018; Torfi et al., 2022; Wu et al., 2020; Navarro et al., 2022; Wagner et al., 2023; Herbach, 2021; Si & Roberts, 2021; Narayanan et al., 2022; Rayhan et al., 2017; Yang et al., 2022a, b; Harer et al., 2018; Lee et al., 2020; Zavrtanik et al., 2021; Rezvani et al., 2021; Prapas et al., 2023; Thambawita et al., 2020; Vijayan et al., 2017; Brophy & Lowd, 2020; Lyu et al., 2021; Chakraborty et al., 2023; Rajabi & He, 2021; Kim et al., 2022; Kiran et al., 2018; Mousavian et al., 2016; Rohani & Eslahchi, 2019; Rao et al., 2022). These papers were published in venues ranging from arXiv only to Nature Scientific Reports, ICCV, ECCV, and Bioinformatics, among others. This reflects not only the widespread belief in this claim, but also that we may be too comfortable making seemingly “correct” assertions without appropriate attribution in ML today.
This claim is frequently attributed to papers that do not make this claim.
Among the 97 that reference this claim and cite a source for this assertion, 39 do not cite any papers that make this claim in the first place (Yang et al., 2015; Li et al., 2020; Kyono et al., 2018; Seo et al., 2021; Hong et al., 2019; Hagedoorn & Spanakis, 2017; Babaei et al., 2021; Zou et al., 2022; Mangolin et al., 2022; Mosteiro et al., 2021; Showalter & Wu, 2019; Cranmer & Desmarais, 2016; Bryan & Moriano, 2023; Zhang et al., 2017; Domingues et al., 2020; Shukla & Marlin, 2019; Blevins et al., 2021; Hsu et al., 2020; Smith et al., 2023; Chu et al., 2018; Deshwar et al., 2015; Mongia et al., 2021; Rubin et al., 2012; Ahmed & Courville, 2020; Gong et al., 2021; Shukla & Marlin, 2018; Ma et al., 2022; Lei Ba et al., 2015; Newby et al., 2022; Ando & Huang, 2017; Stolman et al., 2022; Won et al., 2019; Stephenson et al., 2022; Srivastava et al., 2019; Karadzhov et al., 2022; Vens et al., 2008; López et al., 2013; Hall et al., 2023; Goyal & Khiari, 2020). In total, 13 sources are cited that neither reference nor argue this claim (Davis & Goadrich, 2006; Branco et al., 2016; Provost & Fawcett, 1997; Sokolova & Lapalme, 2009; Wahid-Ul-Ashraf et al., 2019; Ezzat et al., 2017; Burez & Van den Poel, 2009; Flach et al., 2011; Krawczyk, 2016; He & Garcia, 2009; LCT14558, 2017; Lobo et al., 2008). Most often, papers erroneously attribute this claim to Davis & Goadrich (2006), which was cited as a source for this claim 47 times. While Davis & Goadrich (2006) makes many interesting, meaningful claims about the ROC and PR curves, and does argue that the precision-recall curve is more informative than the ROC in cases of class imbalance it never asserts that the area under the PR curve should be preferred over the area under the ROC in cases of class imbalance. It references the emergence of the use of AUPRC instead of or in addition to the AUROC in this context, citing among those references a paper that would later be re-published as (Goadrich et al., 2006), which does make this claim, but (Davis & Goadrich, 2006) itself makes no claim about whether or not AUPRC should be preferred in this way, even by proxy to those prior references. The fact that, despite this, it receives so much citation volume for this claim reflects poorly on the accuracy of our scientific discourse in ML today.
Arguments associated with this claim frequently over-generalize the applicability of this claim or are biased by the appearance of metric superiority rather than true metric utility.
@@ 5 If AUPRC is not better than AUROC under class imbalance, why did we think it was?
As noted in Section 4, there are real world settings in which AUPRC is more aligned with real-world usage than is AUROC (e.g., in a single-group, top-$k$ retrieval setting). However, Claim 1 is often made as a statement about all settings featuring class imbalance. Given this over-generalization, it is unsurprising that many arguments presented in favor Claim 1 when it is used in the literature are similarly over-generalized beyond the cases in which they would be appropriate. For example, claims such as that “precision-recall curves are more informative of deployment metrics” are often used to justify why AUPRC should be used in all cases of class imbalance, rather than just in cases where the relevant deployment metrics are most directly associated with the PR curve. Another class of arguments made in favor of Claim 1 can be reduced to arguments that the metric AUROC is poor in cases of class imbalance because the scores it produces are misleadingly high. While this argument can reflect a meaningful limitation of the communication value of the AUROC, comments about singleton metric results (rather than model comparison through metric values) are inherently orthogonal to the goal of model evaluation. In other words, what matters for model evaluation is not how high a given metric is, but rather the extent to which the metric meaningfully captures the right improvements in the model in the right ways. For a full breakdown of the arguments we observed in the literature and the sources making them, see Appendix Tables 2 and 3.
5.2 Finding 2: Claim 1 has led to inappropriate evaluation usage across a variety of high impact settings and in high-impact publication venues
We find that Claim 1 has been used to justify the use of AUPRC in a variety of settings where AUROC would actually be a superior evaluation metric, in high-impact domains including healthcare and safety, across high-impact venues including NeurIPS, AAAI, Cancer Cell, Nature Scientific Reports, Briefings in Bioinformatics, and Critical Care Medicine, among others. Many of the papers in question here have been cited numerous times, further underscoring the potential negative impact erroneous metric choices in these sources could have. These sources include the following works: (Wagner et al., 2023; Yuan et al., 2015; Lim & van der Schaar, 2018; Leisman, 2018; Cho et al., 2021; Kyono et al., 2018; Yang et al., 2022b; Meister et al., 2022; Mosteiro et al., 2021; Hashemi et al., 2018; Ozyegen et al., 2022; Thambawita et al., 2020; Hsu et al., 2020; Choi et al., 2018; Tiulpin et al., 2019; Lopez-Martinez et al., 2022; Gong et al., 2021; Moor et al., 2019; Ding et al., 2018).
For a representative example of these works, consider Wagner et al. (2023), in which the authors use ML methods to predict colorectal cancer status from medical imaging data. They suggest the following workflow for their model “Our intended clinical use of this workflow is as follows… First, a patient attends a clinic either with suspected CRC or for routine CRC screening. A colonoscopy shows a suspicious tumor, which is evaluated histologically and found to be an adenocarcinoma… Because of its high sensitivity, our algorithm could serve as a filtering step followed by affirmative testing for MSI-high predicted cases. Applying AI-based biomarker prediction would reduce the additional testing burden and therefore speed up the step between taking the biopsy and the molecular determination of MSI-high status, thus enabling an earlier treatment with immunotherapy if indicated.” Under this workflow, two things are clear: (1) the cost of a false positive in this setting is comparatively lower than a false negative, as a false positive only results in unnecessary “molecular determination of MSI-high status” whereas a false negative results in a delay of treatment with immunotherapy, and (2) as motivated by this cost ratio, the model is appropriately designed for (and achieves) a high sensitivity (a.k.a. recall).
As illustrated visually in Figure 1 and argued from first principles through our theoretical analyses in Sections 2 and 4, in this setting given points 1 and 2 above, AUPRC is very much not the right metric to use, precisely because it will favor improvements in high-score regions that will optimize precision rather than those in low-score regions that will help minimize false negatives. Despite this, the authors of this work, as motivated by Claim 1’s extensive publication history, explicitly use AUPRC over AUROC in this setting due to the class imbalance of their problem. This demonstrates the real-world negative impact that the spread of Claim 1 has had in the scientific community.
@@ 6 Limitations and Future Works
While our analyses are thorough and compelling, there are still a number of areas for further improvement and future work. Firstly, our theoretical findings can be refined and generalized to less restrictive settings, that take into account the difficulty of the target task (which may differ between subgroups) or does not require models to be calibrated (in the case of Theorem 3). Further, extending our real-world experiments to more fairness datasets and identifying more nuanced ways to probe the impact of metric choice on disparity measures would significantly strengthen this work. Lastly, These analyses can be extended to consider other metrics, such as the area under the precision-recall-gain curve (Flach & Kull, 2015), the area under the net benefit curve (Talluri & Shete, 2016; Pfohl et al., 2022), and single-threshold, deployment centric metrics as well. In addition, further expanding theoretical analyses, such as examining how Theorem 3 can be softened to have less restrictive requirements regarding calibration or whether or not these results imply anything about the viability or safety of post-hoc calibration of models optimized either through AUPRC or AUROC would also be extremely interesting.

### v4

@@ Results.
Our results demonstrate the impact of the optimization metric on subpopulation disparity. In particular, in Figure 2, we observe a notable disparity introduced when optimizing under the AUPRC metric regardless of the optimization procedure. This is evident in the performance metrics across the high and low prevalence subpopulations, which exhibit significant divergence as the optimization process favors the group with higher prevalence. In the more realistic, random-permutation optimization procedure (Figure 2d), this even results in a decrease in the AUROC for the low prevalence subgroup. In comparison, when optimizing for overall AUROC, the AUROCs of both groups increase together. Note that we show the effect of this optimization on the AUPRC metric, which shows very similar trends, in Appendix Figure 4. These results demonstrate explicitly that not only does optimizing for AUPRC differ greatly than for AUROC, as has been noted historically by researchers developing explicit AUPRC optimization schemes [409], but it in fact does so in an explicitly discriminatory way in realistic scenarios.
@@ 3.2 Real-world experimental validation
To demonstrate the generalizability of our findings to the real world, we evaluate fairness gaps induced by AUROC and AUPRC selection on four common datasets in the fairness literature [441, 99, 205].
@@
We use the following four tabular binary classification datasets: adult [17], compas [14], lsac [413], and mimic [178]. In each dataset, we consider both sex and race as sensitive attributes. To mimic the setting of our theorems, we balance each dataset by the sensitive attribute during both training and test, by randomly subsampling the majority group. Further details about each dataset, as well as preprocessing steps, can be found in Appendix H.
@@
We train XGBoost models [65] on each dataset. For each task, we iterate over a grid of per-group weights in order to create a diverse set of models that favor different groups. For each setting of task and per-group weight, we conduct a random hyperparameter search [37] with 50 runs. Though more complex hyperparameter search methods such as BOHB [100] or TPE [36] might lead to better performance, random searches are far more popular and practical, and have been used in popular benchmarking libraries [131, 337].
We evaluate the validation set overall AUROC and AUPRC. We also evaluate the test set AUROC gap and AUPRC gap between groups, where gaps are defined as the value of the metric for the higher prevalence group minus the value for the lower prevalence group. Based on our theorems, our hypothesis is that overall AUPRC should be more positively correlated with the signed AUROC gap than overall AUROC, indicating that it better favors the higher prevalence group, especially when the prevalence ratio between groups is high. To test this hypothesis, we evaluate the Spearman correlation coefficient between these quantities. We repeat this experiment 20 times, with different random data splits, to obtain a 95% confidence interval.
Figure 3: Difference in the Spearman’s $\rho$ between the test-set signed AUROC gap versus the validation set overall AUPRC, and the AUROC gap versus the overall AUROC. Numbers in parentheses are the prevalence ratios between the two groups for the particular attribute, and datasets are sorted by this quantity. Error bars are 95% confidence intervals from 20 different random data splits.
@@ Results.
In Figure 3, we plot the difference in the Spearman correlation coefficient of the AUROC gap versus the overall AUPRC, and AUROC gap versus overall AUROC. We observe mixed results in datasets with low prevalence ratio. In dataset with higher prevalence ratio, we find that overall AUPRC is more positively correlated with the AUROC gap than overall AUROC, indicating that AUPRC more aggressively favors the higher prevalence group. We emphasize that the prevalence ratios observed in these real-world datasets is much lower than the ratio of 5 used in our synthetic experiments, which may account for the mild effect observed. To see raw results from these experiments, see Appendix Figure 7. Similar results for neural network classifiers can be found in Appendix Section H.3.
Next, in Appendix Figure 8, we plot the difference in the Spearman’s $\rho$ from Figure 3, versus the prevalence gap. We find that there is a statistically significant correlation between the two (Spearman’s $\rho$ = 0.905, p = 0.002). Thus, while our power to detect a prevalence mediated AUPRC bias amplification effect is limited due to the limited prevalence disparities in these datasets, we nonetheless observe a strong positive correlation between the extent of the prevalence mismatch between the low and high prevalence group and the amount that AUPRC favors the high prevalence group over AUROC. In other words, our results show that across these fairness datasets and attributes, as the prevalence disparity grows more extreme, we observe a statistically significant corresponding increase in the extent to which AUPRC introduces algorithmic bias, exactly in accordance with what Theorem 3 suggests.
4 When Should One Use AUPRC vs. AUROC?
In Sections 2 and 3, we have shown that AUPRC is not universally superior in cases of class imbalance (and that instead, it merely preferentially optimizes high-score regions over low-score regions) and that it also poses serious risks to the model fairness in settings where subgroup prevalences differ. In light of this, how should we revise Claim 1 to reflect when we actually should use AUPRC instead of AUROC or vice versa? Below, we explore this question and provide practical guidance on metric selection for binary classification, building on our theoretical results and highlight specific contexts in which one may be favorable (Figure 1). Note that while we provide guidance below on situations in which AUROC vs. AUPRC is more or less favorable, this is not to suggest that authors should not report both metrics, or even larger sets of metrics or more nuanced analyses such as ROC or PR curves; rather this section is intended to offer guidance on what metrics should be seen as more or less appropriate for use in things like model selection, hyperparameter tuning, or being highlighted as the ‘critical’ metric in a given problem scenario.
@@
For model evaluations conducted outside of specific deployment contexts, where the differential costs of errors are undefined, the necessity for a metric that impartially values improvements across the entire model output space becomes paramount (Figure 1a). As it is not known in advance where samples of interest will live in the output space, nor are particular cost ratios known, there should be no preference for mistake correction. Therefore, in this setting, AUROC is favorable as it uniformly accounts for every correction, offering a fair assessment irrespective of decision thresholds.
@@
In applications where the consequences of false negatives are especially high, such as in the early screening for critical illnesses like cancer (Figure 1c), a primary goal of the model will be to achieve the fewest missed cancer cases; equating to prioritizing model recall. In such a scenario, the most important mistakes to correct occur at lower score thresholds, as high-score mistakes will not change which positive samples are missed in deployment settings (as chosen thresholds are likely to be low). This behavior is the inverse of what AUPRC prioritizes, demonstrating that in such situations, AUROC should be preferred over AUPRC.
For ethical resource distribution among diverse populations, use AUROC:
When faced with the challenge of ethically allocating scarce resources across a broad population, necessitating equitable benefit distribution among subgroups (Figure 1d), one must avoid prioritizing model improvements that selectively favor one subpopulation. As AUPRC will target high-score regions selectively, it risks unduely favoring high-prevalence subpopulations, as shown in Theorem 3 and Figures 2 and 3. Even though in this resource distribution problem, high-score regions are selectively important compared to low-score regions, the fact that in this problem, we must prioritize across all subpopulations equally means that AUPRC’s global preference is untenable as it could induce bias.
@@
5 Literature Review: Examining how Claim 1 Became so Widespread
Claim 1 states that “AUPRC is better than AUROC in cases of class imbalance” and is widespread in the literature. Via both a manual literature search and an automated search of over 1.5M arXiv papers (see Appendix I for methodology), we observed 424 publications making this claim.333 Note that throughout this section, full citations for all of the larger lists of papers we reference will be relegated to Appendix Section I for concision. This widespread disemmination of Claim 1 has a significant impact on the validity of scientific discourse. We observed examples of high-profile papers operating in medically critical settings where high-recall is a key priority evaluating ML systems via AUPRC due to their task’s underlying class imbalance [399]; papers focusing on fairness critical applications relying on AUPRC due to this claim, even while our results demonstrate AUPRC has major problems in the fairness regime [366, 306], and numerous other papers perpetuating this source of scientific misinformation.
Among the 424 papers we discovered referencing this claim, 167 did so with no associated citation. These papers were published in a wide range of venues, including high profile venues such as NeurIPS, ICML, and ICLR. This reflects not only the widespread belief in this claim, but also that we may be too comfortable making seemingly “correct” assertions without appropriate attribution in ML today. Further, Among the 257 that reference this claim and cite a source for this assertion, 135 do not cite any papers that actually make this claim in the first place. Most often, papers erroneously attribute this claim to [83], which was cited as a source for this claim 144 times. While [83] makes many interesting, meaningful claims about the ROC and PR curves, and does argue that the precision-recall curve may be more informative than the ROC in cases of class imbalance it never asserts that the area under the PR curve should be preferred over the area under the ROC in cases of class imbalance. This distinction is important, because while curves can be used to simultaneously communicate many different performance statistics to their viewers across different FPR/TPR or Precision/Recall criteria, and therefore should be assessed primarily as communication tools to deployment experts, areas under these curves are single-point summarizations which are primarily used for deployment-agnostic model comparison, which is a very different use case.
@@ 5 Literature Review: Examining how Claim 1 Became so Widespread
Even when appropriate papers are cited, the valid settings in which AUPRC should be preferred (see Section 4 for examples) are often over-shadowed by significant over-generalizations to preferring AUPRC in all settings featuring class imbalance. For example, claims such as that “precision-recall curves are more informative of deployment metrics” are often used to justify why AUPRC should be used in all cases of class imbalance, rather than just in cases where the relevant deployment metrics are most directly associated with the PR curve. Another class of arguments made in favor of Claim 1 is rooted in claims that AUROC is poor in cases of class imbalance because its scores are misleadingly high. While this argument can reflect a meaningful limitation of the communication value of the ROC or the AUROC, comments about singleton metric results (rather than model comparison through metric values) are inherently orthogonal to the goal of model evaluation. In other words, what matters for model evaluation is not how high a given metric is, but rather the extent to which the metric meaningfully captures the right improvements in the model in the right ways. The widespread nature of Claim 1 has also led researchers astray when exploring new optimization procedures for AUPRC, by advocating for the importance of AUPRC when processing skewed data, even in domains such as medical diagnoses that often have high false negative costs relative to false positive costs [409]. For a more extensive breakdown of the arguments we observed in the literature and the sources making them, see Appendix Tables 2 and 3.
@@ 6 Limitations and Future Works
There are still a number of areas for further improvement and future work. Firstly, our theoretical findings can be refined and generalized to, e.g., take into account the difficulty of the target task (which may differ between subgroups), not require models to be calibrated (in the case of Theorem 3), or specifically take into account more than 2 subpopulations for more nuanced comparisons beyond what can be inferred through pairwise comparisons between subpopulations, where our results would naturally apply. Further, extending our real-world experiments to more fairness datasets and identifying more nuanced ways to probe the impact of metric choice on disparity measures would significantly strengthen this work. These analyses can also be extended to consider other metrics, such as the area under the precision-recall-gain curve [104], the area under the net benefit curve [384, 307], and single-threshold, deployment centric metrics as well. In addition, one of the largest limitations of Theorem 3 is its restrictive assumptions, in particular the requirement of perfect calibration. A ripe area of future work is thus to investigate how we can soften our analyses for models with imperfect calibration or to determine whether or not our results imply anything about the viability or safety of post-hoc calibration of models optimized either through AUPRC or AUROC.